Solve the following quadratic equations by using an optimum method.
x^2 - 9x - 4 = 6
Please help.
Thank you.

Found 2 solutions by stanbon, Wahmed301@:
Answer by stanbon(75887)

Subtract "6" from both sides of the equation to get:
x^2-9x-10=0
Factor to get:
(x-10)(x+1)=0
Then either x-10=0 or x+1=0
So, either x=10 or x=-1
Cheers,
Stan H.

Answer by Wahmed301@(1)